Problem 36.39 Part A At what speed, as a fraction of c, is a particles kinetic energy twice its rest energy Express your answer using three significant figures Submit My Answers Give Up
Answer #1

the relation between kinetic energy and rest energy is

KE = (\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}-1) * E

here it is given KE = 2 * E

so (\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}-1) = 2

\frac{1}{\sqrt{1-\frac{v^{2}}{c^{2}}}}=3

1-\frac{v^{2}}{c^{2}}=\frac{1}{9}

v = 0.9428 c